INAF queue observing, 2018 May 08-09

Observer: F. Cusano, A. Gargiulo
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument:MODS, LUCI

Summary

At the sunset sky is clear
instruments: MODS2 red grating is not available all the other instruments are fine
Night duration (12\xB0 el twilight) 3:30- 11:00 UT = 7 h 30m

Observing time: 5h 40m (75%)

Weather time loss: 0h 30 (7%)

technical time loss 1h 20m (18%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
sky flat   LBC-R Y, i-sloan      
Gilli Mask 22 MODS1+LBC-R dual grating, i- sloan 2/4 h mods1r.20180509.0004-9, mods1b.20180509.0001-0006 1.3h of i images
Standard Feige 34 MODS1 dual grating, red grating      
Severgnini J1431 LUCI2 AO - K band 0.5 /0.6 h   target completed FWHM ~0.15", no psf star, check log
Prandoni USS202 LUCI1+2 Ks, J,z 0.9/0.9h, 1/1h, /1.7h   z images wih clouds, to be checked
Mods Call mask22 MODS1 dual grating     arc, flat Gilli
Luci cal flat LUCI2 AO - K band     Severgnini cal
LBC cal bino bias LBC-B-R       bias
Luci cal flat LUCI1+2 Ks, J, z     Prandoni cal
Mods Call ls MODS1 dual grating, red grating     arc, flat lfor s and standard

Details

2:16 starting sky flat with LBC-R, Y and i-sloan

2:53 slewing to Gilli mask 22
  • 2:55 dohybrid LBC-R, not converging at 3:08
  • acq mods1r.20180509.0001-0003
  • 3:06 spec mods1r.20180509.0004-0009, mods1b.20180509.0001-0006
  • 3:07 seeing 0.8" on guider
  • 3:16 LBC-R converged, starting i-sloan imaging, i_sloan.xml
  • 3:30 seeing 0.9" on guider
  • 3:53 stop LBC obs stars elongated, 3:54 running dohybrid
  • 4:03 converged, starting again i_sloan.xml
  • 4:05 seeing 0.9" on guider
  • 4:28 seeing 0.9" on guider
  • 4:49 stop mods script at exp 5 of 6 to power off dx-sex, some problem that can affect the alert system, the rest is fine
  • 5:00 stop also LBC stars elongated
  • 5:03 starting again the mods script
  • 5:03 dofpia
  • 5:06 seeing 1.1" on guider

5:30 slewing to standard Feige 34
  • acq mods1r.20180509.0010-0011
  • spec dual grating mods1r.20180509.0012-0014, mods1b.20180509.0007-0009
  • spec red mods1r.20180509.0015-0017

06:01 switch to LUCi to perform AO program

06:22 slewing to Severgnini program, target J1431

06:30 problem with the right secondary not converging, not working we use LUCI1 only

06:37 sending script SDSSJ1431_sci.xml, present cancelled, authorizing LUCI1 -only

06:44 right side working again we send the binocular script again, sending binocular obs

07:11 we observe with LUCI2 only, executing script SDSSJ1431_sciluci2.xml, measured FWHM 0.15"

07:48 seeing 0.7" on guider, 1" dimm

07:57 starting psf star of Servergnini, AO star to faint, but is 15.3 from catalog
  • star is to faint we do not observe the psf sfar, otherwise in the scientific field of view there are two stars that can be used to reconstruct the psf

08:10 slewing to Prandoni K USS202,
  • binocular observations uss202_Ks_bino.xml
  • seeing
  • 8:13 check pointing cancelled present and mask problem on luci2

8:48 mask problems solved, we send again the script Ks ofr USS202
  • 8:49 dimm seeing 0.7', seeing 0.7" on guider
  • 9:25 seeing 0.7" on guider

9:31 sending USS202 J script
  • 9:05 1.2" seeing on guider at 10:02, dimm seeing 1"
  • 9:10 some problem of collimation on the left side (LUCI1)
  • 10:11 clouds coming in, guide star still good, clouds did not affected the ob that ended at 10:15

10:16 guide star is not fading, the clouds are not on the target, we send the script for the z band USS202
  • 10:20 seeing 0.8" on guider
  • 10:22 guide star dimming of 0.2 mag
  • 10:24 script paused for clouds
  • 10:33 script resumed the guide star is back to its magnitude
  • 10:53 clouds are gone seeing 0.7", data are good from this point

11:44 mods cal for Gilli mask22
  • flat mods1r.20180509.0018-0023, mods1b.20180509.0010-0015
  • arc mods1r.20180509.0024-0026, mods1b.20180509.0016-0018

11:40 starting LUCI cal
  • LUCI flat for Severgnini program, N30, K filter images luci2.20180509.0098-0108, the mask is not perfectly aligned as it was in the observations, * the SA moves the mask manually in the same position and we take again flats luci2.20180509.0115-124 use this flats

  • MODS grlamp mods1r.20180509.0027-0029, mods1b.20180509.0019-0021
  • sliflats 5 arcsec mods1r.20180509.0030-0032, mods1b.20180509.0022-0027
  • grlapms_r mods1r.20180509.0033-35
  • slitflats_r for Caccianiga slit 1.2" mods1r.20180509.0036-40
  • sliflats 5 arcsec red mods1r.20180509.0041-44
  • bias 8 k mods1r.20180509.0045-0049, mods1b.20180509.0028-0032
  • bias 3k mods1r.20180509.0050-0054, mods1b.20180509.0033-0037
  • grpixflats mods1r.20180509.0055-59, mods1b.20180509.0038-0047
  • grpixflats_r mods1r.20180509.0060-0064
  • LBC bino bias

  • Prandoni dome flat Ks, J, z

 
Topic revision: r7 - 04 Jul 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback